OpenStudy (anonymous):

Elena joined a golf club two years ago. She pays an annual membership fee of $925 and a greens fee of $38 each time she plays a game of golf. The function below can be used to calculate her total yearly golfing fee, f (g), in dollars. f (g) = 925 + 38g , where g represents the number of times she plays golf during the year. Last year she paid $1,913 as a total golfing fee. For how many games did she pay a greens fee? The answer is 26 right? 1913 is f(g) solve for g (1913 - 925)/38 = 26

OpenStudy (anonymous):

Crew's surf shop charges $15 per hour to rent paddle boards and $5 per hour to rent a wetsuit. Kai and Reyna paid a total of $260 to rent a paddle board and a wetsuit each one day. Reyna rented them for 3 hours less than Kai. The following equation represents this situation, where x represents the number of hours Kai rented a surfboard and a wetsuit. 15x + 5x + 15(x − 3) + 5(x − 3) = 260 What is the total amount Kai paid to rent the paddle board and wetsuit that day? I got 160 for my answer, is that correct?

OpenStudy (anonymous):

A website recorded 1,351 visitors one month. This was 30% less than the number of visitors to the site the previous month. The following equation represents this situation, where x represents the number of visitors to the site the previous month. x − 0.3x = 1,351 How many visitors did the website record in both months combined? is the answer 3108? im not sure

OpenStudy (anonymous):

or maybe it 1930?

OpenStudy (anonymous):

there's a lot I know, im sorry but word problems seem to be the hardest for me, so I want to make sure there right! :)

OpenStudy (anonymous):

No problem. The first one is right. The last one is 0.7x = 1351 => x = 1351/0.7 = 1930. 1930 + 1351 = 3281

OpenStudy (anonymous):

and what about the one in the middle there?

OpenStudy (anonymous):

Seems correct.

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

okay, cool would you mind me going to yahoo answers, writing the questions i need checked and the answers and then send you the link for it, it mite be easier then having to write them all out on here? if you have the time that is...

OpenStudy (anonymous):

I am a bit busy right now, maybe if you post it one at a time, people can answer for you :-) Good luck with it :-)

OpenStudy (anonymous):

okay, thanks :)